The product of two divergent sequences is divergent. True or false? Justify.


1
Expert's answer
2021-04-14T14:29:02-0400

The product of two divergent sequences is divergent "\implies" it is a False statement

Counter example:

Let two sequences: X = (1, 0, 2, 0, 3, 0, ............. )

Y = (0, 1, 0, 2, 0, 3, ..............)

clearly, both sequences are unbounded and diverging to "\infty"

there product X"\cdot"Y = (0, 0, 0, 0, 0, 0, .........)

which is a constant sequence converging to 0

therefore,

X"\cdot"Y is convergent sequence.